OpenStudy (anonymous):

Two lines are parallel. One passes through the points (-3,4) and (5,k) and the other has slope -2. What is the value of k?

OpenStudy (anonymous):

y = mx + b 4 = -2(-3)+b 4 = 6 + b -2 = b y = -2x -2 equation...no substitute 5 in for x, and you'll get k

OpenStudy (anonymous):

k = -2(5)-2 k = -10-2 k = -12

OpenStudy (anonymous):

Another way...y1 - y2 / x1-x2 4-k / -3-5 = -2 4-k/-8 = -2 4-k = 16 -k = 12 k=-12